Will all the lead be removed from 12.2 mL of 7.00×10-3 M Pb(NO3)2 upon addition of 12.5 mL of 0.0115 M Na2S? If all the lead is removed, how many moles of lead is this? If not, how many moles of Pb remain?
1
Expert's answer
2017-10-12T02:52:07-0400
12.2 mL of 7.00×10-3 M Pb(NO3)2
12.5 mL of 0.0115 M Na2S
Pb(NO3)2 + Na2S = PbS + 2NaNO3
1) n(Pb(NO3)2) = 12.2*10^-3 * 7.00*10^-3 = 85.4*10^-6 mol
2) n(Na2S) = 12.5*10^-3 * 0.0115 = 143.8*10^-6 mol
all the lead is removed n(Pb) = 85.4*10^-6 mol